OpenStudy (anonymous):

can anybody help me with this A farmer needs to mix 12 ounces of fertilizer for every 10 pounds of soil. How many pounds of fertilizer will the farmer need to make a mix that contains 120 pounds of soil?

OpenStudy (anonymous):

\[(120/10) * 12 = y\]

OpenStudy (anonymous):

\[\frac{ 12 }{ 10 } = \frac{ x }{ 120 }\] where x is the yes

OpenStudy (anonymous):

12 * 120 is equal to what?

OpenStudy (cheesecakekitten):

gOOD JOB, EVERYONE

OpenStudy (anonymous):

Since\[120/12 = 10\] Then your equation is \[12 * 12 = y\] Which one is it?